A wire of length y meters is bent to form a square of perimeter y meters. If Area of square is twice its perimeter, then find the length of wire
Answer by Alan3354(31514)

The sides will be y/4
Area = (y/4)^2 = y^2/16
y^2/16 = y (numerically equal)
y^2 = 16y
--> y = 0
y = 16